Quatro Co. issues bonds dated January 1, 2019, with a par value of $790,000. The bonds annual contract rate is 9%, and interRequired 1 Required 2 Required 3 How much total bond interest expense will be recognized over the life of these bonds? TotalRequired 1 Required 2 Required 3 Prepare an effective interest amortization table for these bonds. (Round all amounts to the
Answer #1
1
Premium 20694 =810694-790000
2
Total interest expense over life of bonds
6 payments of $ 35550 213300
Par value at maturity 790000
Total repaid 1003300
Less: Amount borrowed 810694
Total bond interest expense 192606
3
Semiannual Interest period end Cash interest paid Bond interest expense Premium amortization Unamortized Premium Carrying value
01/01/2019 20694 810694
06/30/2019 35550 32428 3122 17572 807572
12/31/2019 35550 32303 3247 14325 804325
06/30/2020 35550 32173 3377 10948 800948
12/31/2020 35550 32038 3512 7436 797436
06/30/2021 35550 31897 3653 3783 793783
12/31/2021 35550 31767 3783 0 790000
Total 213300 192606 20694
Workings:
Cash interest paid 35550 =790000*9%*6/12
Bond interest expense = Carrying value X 8% X 6/12
Bond interest expense:
06/30/2019 32428 =810694*8%*6/12
12/31/2019 32303 =807572*8%*6/12
06/30/2020 32173 =804325*8%*6/12
12/31/2020 32038 =800948*8%*6/12
06/30/2021 31897 =797436*8%*6/12
12/31/2021 31767 =793783*8%*6/12